OpenStudy (anonymous):

Medals Medals Medals! im trying to find the exact value of each equation, a little help? 1. cot[cot^-1sqrt3/3] 2. sin^-1(cos pi/2) 3. tan(sin^-1x/2)

OpenStudy (anonymous):

2. \[-\frac{\pi }{6} \]3.\[\frac{x}{2 \sqrt{1-\frac{x^2}{4}}} \]

OpenStudy (anonymous):

okay thank you! i can see that. Do you know anything for one?

OpenStudy (anonymous):

Use the fact that \(\cot(\cot^{-1}x)=x\).

OpenStudy (anonymous):

alright, i recall that! thank you!

OpenStudy (anonymous):

By the way, the second one's not right. \[\cos\frac{\pi}{2}=0\] so \[\sin^{-1}0=0\]

OpenStudy (anonymous):

thanks so much for that, really!

OpenStudy (anonymous):

2. According to Mathematica v9 and WolframAlpha, and with no delimiters: \[\sin ^{-1}\left(\frac{\cos (\pi )}{2}\right)=-\frac{\pi }{6} \]